Assume that GDP is Rs. 6000, personal disposable income is Rs. 5100 & the Govt. Budget Deficit is Rs. 200, consumption is Rs. 3800 & trade deficit Rs. 1000. Calculate saving, investment & government spending?
1
Expert's answer
2017-01-10T09:37:14-0500
Y = Rs. 6000, Yd = Rs. 5100, BD = 200, C = Rs. 3800, NX = -1000.
Saving S = Yd - C = 5100 - 3800 = Rs. 1300,
Y = C + I + G + NX,
Government spending G = T + BD, T = Y - Yd = 6000 - 5100 = 900, so G = 900 + 200 = Rs. 1100.
Investment I = Y - C - G - NX = 6000 - 3800 - 1100 + 1000 = Rs. 2100.